A stream of water gushes from an open pipe at an angle of 15o from the horizontal. If water comes out of the pipe at a speed of 15 m/s, how far from the pipe will the water land?


1
Expert's answer
2021-10-24T18:18:27-0400

The range is given as follows (https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):



"R = \dfrac{u^2\sin (2\theta)}{g}"

where "u = 15m\/s""\theta = 15\degree" and "g = 9.8m\/s^2"

"R = \dfrac{15^2\cdot \sin(2\cdot 15\degree)}{9.8} \approx 11m"

Answer. 11 m.